OpenStudy (anonymous):

solve sinx^2 + cos2x - cosx = 0 @StudyGurl14

OpenStudy (studygurl14):

Sorry, I thought I could help, but know I'm not so sure. @jim_thompson5910

OpenStudy (campbell_st):

use the substitution \[\cos^(2x) = \cos^2(x) - \sin^2(x)\] and things should become easier

OpenStudy (anonymous):

but when you solve equations like that you are supposed to set one side to 0 so you can factor

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

i get it, thanks!! but what about the sine and cosine?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

nevermind, i get it all now thanks!

OpenStudy (campbell_st):

well you have \[\sin^2(x) + \cos^2(x) - \sin^2(x) - \cos(x) = 0\] so what cancels out...? then you can factor what's left for the solutions